OpenStudy (anonymous):

Probabilty: A Motor Vehicle Number Plate Consists Of 3 Letters And Followed By 3 Digits / How Many Different Number Plated Are Possible?

OpenStudy (ash2326):

There are 26 letters A-Z 10 numbers 0-9 It's not mentioned that repetition is not allowed. So we have 5 places to fill, first three for letters other two for numbers \[\large \text{_ _ _ _ _}\] Can you try?

OpenStudy (anonymous):

mmm not really :S

OpenStudy (ash2326):

The first place can be filled in 26 ways. the second and third are also letters, so they would be filled in 26 ways as well. Total no. of ways \[26\times 26\times 26\times -\times -\] Now you need to fill the blanks for no.s , remember you have the digits 0-9

OpenStudy (anonymous):

thank u so much!!!!